B.        The table below lists the sales, y (in millions of dollars) and the number of employees, x (in thousands) for a random sample of 20 Fortune 500 companies. The regression results based on the model are given below. Some of the numbers in the regression tables have been taken out.

SUMMARY OUTPUT

Regression Statistics

Multiple R

(1)

R Square

(2)

Adjusted R Square

0.837364

Standard Error

(3)

Observations

(4)

ANOVA

df

SS

MS

F

Significance F

Regression

(5)

(7)

(9)

(10)

9.7801E-09

Residual

(6)

(8)

7825568.1

Total

19

914226344.8

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

816.7943

767.9534976

1.06359862

0.3015718

-796.617401

2430.206

Employees, x

(11)

15.04320104

9.94110437

9.78E-09

117.9414144

(12)

B1.      Answer for (1):

            a.         0.8374

            b.         9.9411

            c.         0.9197

            d.         816.7943

B2.      Answer for (2):

            a.         0.9197

            b.         0.8459

            c.         0.8374

            d.         20

B3.      Answer for (3):

            a.         98.8256

            b.         0.8459

            c.         2797.4217

            d.         15.0432

B4.      Answer for (4):

            a.         2797.4217

            b.         20

            c.         18

            d.         117.9414

B5.      Answer for (5):

            a.         1

            b.         20

            c.         18

            d.         98.8256

B6.      Answer for (6):

            a.         1

            b.         18

            c.         19

            d.         20
Answer #1

B1)

Multiple R =sqrt(1-SSE/SST) =sqrt(1-7825568.1*18/914226344.8)=0.9197

B2)

R square =0.9197^2 =0.8459

B3)

Standard Error =sqrt(MSE )=sqrt(7825568.1) =2797.4217

B4)

number of observation =total (df)+1 =19+1 =2

B5) df(regression) =1

B6 ) residual df =n-2=20-2 =18